# Dynamics of Moral Behavior in Heterogeneous Populations of Learning Agents

### v3

@@ Appendix A Partner Selection Algorithm
The algorithm for partner selection ($sel$) & at the basis of the dilemma game ($dil$) with 2 separate Q-network models is presented below (see Algorithm 1). We use episodes, so that all players learn simultaneously, after each player had a single chance to select and play. For clarity, each selector’s state $s_{i,sel}^{t}$ relies on the same $environment\_state$ (i.e., simultaneous selections), and therefore every player $i$’s $s_{i,sel}^{t+1}$ in the $sel$ trajectory only differs from $s_{i,sel}^{t}$ by one entry: $a_{j,dil}^{t}$ at player $j$’s position. For $dil$, each player $i$ uses all of their observed dilemma trajectories to update their network - by calculating an average loss across these trajectories.
@@
population of players $P$ of size $N$, total episodes $T=30000$
set random initial $environment\_state$ (i.e., one random move for each player)
for all players $i\in P$ do:
@@
    initialize selection memory $D_{i,sel}$ and dilemma memory $D_{i,dil}$ (capacity $265$ to allow experience replay)
    initialize value functions $Q_{i,sel},Q_{i,dil}$
    set random initial selection & dilemma states $(s_{i,sel}^{t=0},s_{i,dil}^{t=0})$
end for
initialize iterations counter $t=0$
for all ${episode}\leftarrow 1,T$ do:
    Get current global $environment\_state$ from the environment $\triangleright$ this remains stable for entire episode
    PART 1: Partner Selection
    for all players $i\in P$ do $\triangleright$ loop over entire population in a random order
         if ${episode}=0$ then
             use random state stored in $s_{sel}^{t=0}$ on first episode
         else
             $i$ observes the $environment\_state$ and removes itself to parse in their own $s_{i,sel}^{t}$
         end if
         $t\leftarrow t+1$
@@
         using $s_{i,sel}^{t}$, $Q_{i,sel}$ and $\epsilon$-greedy policy, $i$ selects a partner $j$ (i.e., takes action $a_{i,sel}^{t}$)
         store pair of agents $\{i,j\}$ in this episode’s sub-population $P_{episode}$
    end for
    PART 2: Dilemma Game
    for all $\{i,j\}\in P_{episode}$ do $\triangleright$ for each pair of agents resulting from the selection stage
         if $episode=0$ then
             use random dilemma state stored in $s_{dil}^{t=0}$ on first episode
         else
@@
             $i$ and $j$ parse states $s_{i,dil}^{t}$, $s_{j,dil}^{t}$ from the $environment\_state$ the current opponent’s index
         end if
         using $s_{i,dil}^{t}$, $Q_{i,dil}$ and $\epsilon$-greedy policy, $i$ plays a move $a_{i,dil}^{t}$
         using $s_{j,dil}^{t}$, $Q_{j,dil}$ and $\epsilon$-greedy policy, $j$ simultaneously plays a move $a_{j,dil}^{t}$
         $i$ and $j$ receive rewards $r_{i,dil}^{t}$ $r_{j,dil}^{t}$ respectively
         $i$ and $j$ observe $s_{dil}^{t+1}$ based on the current opponent’s action $a_{t}$: $s_{i,dil}^{t+1}=a_{j,dil}^{t};s_{j,dil}^{t+1}=a_{i,dil}^{t}$
         player $i$ copies the dilemma reward to the selection reward as well: $r_{i,sel}^{t}=r_{i,dil}^{t}$
         in player $i$’s $D_{i,dil}$ memory, store: $(s_{i,dil}^{t},a_{i,dil}^{t},r_{i,dil}^{t},s_{i,dil}^{t+1})$
         in player $j$’s $D_{j,dil}$ memory, store: $(s_{j,dil}^{t},a_{j,dil}^{t},r_{j,dil}^{t},s_{j,dil}^{t+1})$triangleright i.e., opponent $j$ learns too
    end for
    PART 3: Update $sel$ and $dil$ networks $\triangleright$ after each player made a selection & played a game:
    for all players $i\in P$ do:
@@
         update the corresponding part of $environment\_state$ with the latest moves by $i$ ($a_{i,dil}^{t}$)
@@
         player $i$ observes their $s_{i,sel}^{t+1}$ by copying the latest $environment\_state$ and removing self
         in player $i$’s $D_{i,sel}$ memory, store: $(s_{i,sel}^{t},a_{i,sel}^{t},r_{i,sel}^{t},s_{i,sel}^{t+1})$
         player $i$ samples one experience from $D_{i,sel}$: $(s_{i,sel}^{t},a_{i,sel}^{t},r_{i,sel}^{t},s_{i,sel}^{t+1})$ calculates loss
         player $i$ samples all experiences from $D_{i,dil}$: $(s_{i,dil}^{t},a_{i,dil}^{t},r_{i,dil}^{t},s_{i,dil}^{t+1})$ calculates average loss
@@
         player $i$ updates $Q_{i,sel}$ network with respect to time step $t$ the calculated loss
@@
         player $i$ updates $Q_{i,dil}$ network with respect to time step $t$ the calculated loss
         refresh memory buffers $D_{i,sel}$, $D_{i,dil}$, $D_{j,dil}$
    end for
end for
Appendix B Players’ Popularity
Figure 5 shows the confidence intervals around the means presented in the main body of the paper for the most popular players in each population on the final 100 episodes (compare with the heat map in Figure 3(a)), and provides the 50% threshold line for ease of interpretation, which allows us to compare whether the majority player is selected more or less often than expected simply due to their prevalence.
Figure 5: Most popular players in each population on the final 100 episodes, with errors bars around the mean 50% reference line. We average across 20 runs.
Appendix C Action Pairs within Each Population
An analysis of action pairs in Figure 6 provides further insight into the pair-wise interactions in each population (including mutual cooperation - C,C; unilateral exploitation - C,D or D,C; and mutual defection - D,D observed at every episode). We observe that the high levels of cooperation in the majority-Ut and majority-V-Ki populations are largely driven by mutual cooperation, and that the low cooperation in the majority-aUt, majority-V-Ag and the majority-V-Eq populations can be largely attributed to high mutual defection. In other populations much more exploitation (i.e., unilateral defection) is observed, and in certain pro-social populations such as majority-De this is largely due to the selecting player being exploited by their selected opponent.
Figure 6: Action pairs observed on every episode in every population (we sum observations within an episode, and then count the number of occurrences of each action pair across 20 runs). The ordering of actions within each pair corresponds to the selecting player first, and the selected opponent second.
Appendix D Cooperation by Each Player Type across Populations
In addition to the analysis of the S agents’ behavior across populations presented in the main body of the paper, Figure 7 presents cooperation over time by every other player type in every population.
Figure 7: Behavior of each player type across every population. We present moving average of the mean across 20 runs.
For all players except V-Eq and V-In, levels of cooperation are relatively consistent across populations. V-Ki and V-Ag players go through an initial phase of low or high cooperation, respectively. The other norm-based pro-social player De converges to full cooperation very quickly, showing that this definition of the norm produces the strongest signal for cooperation in the IPD (extending the findings of Tennant et al. 2023a to the population case as well). Their anti-social counterpart mDe, however, does not display full defection, likely due to the partner selection mechanism providing an incentive to cooperate in order to get selected more often to play. The Ut and aUt players converge to either full cooperation or full defection respectively in all populations nearly equally, but the rate of convergence is faster for the anti-social player, likely due to the payoffs associated with defection in the IPD game. The V-Eq and V-In players cooperate to variable extents in different populations, but in a way that is consistent with the choice of their reward functions (e.g., defect more in a majority-defective population).
Appendix E Selection Patterns between Player Types
In addition to the analysis of selections made over the entire simulation in the two populations presented in the main body of the paper (majority-S and majority-De, see Figure 3(b)&3(c)), we show selections made by each individual player in every population in Figures 8 (as a heat map) and 9 (as networks involving the top 10% of the pairs in terms of number of interactions). The latter show the most common patterns of interaction among players in these artificial societies - for example, in the majority-De population, the De players select other types of opponents more often than each other.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 8: Selection patterns across player types in every population, with number of selections in every cell summed over all 30000 episodes. We average across 20 runs
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 9: Selection patterns across player types in every population (most common interactions, i.e., >90th percentile, over all 30000 episodes). More popular players are placed more centrally using a force-directed algorithm.

### v4

@@ Appendix D Partner Selection Algorithm
The algorithm for partner selection ($sel$) and at the basis of the dilemma game ($dil$) with 2 separate Q-network models is presented below (see Algorithm 1). We use episodes, so that all players learn simultaneously, after each player had a single chance to select and play. For clarity, each selector’s state $s_{i,sel}^{t}$ relies on the same $environment\_state$ (i.e., simultaneous selections), and therefore every player $i$’s $s_{i,sel}^{t+1}$ in the $sel$ trajectory only differs from $s_{i,sel}^{t}$ by one entry: $a_{j,dil}^{t}$ at player $j$’s position. For $dil$, each player $i$ uses all of their observed dilemma trajectories to update their network, by calculating an average loss across these trajectories.
We use the DQN algorithm for agents learning to select and to play the dilemma (Mnih et al. 2015), with a slight simplification: we refresh the experience buffer at the end of every episode (see Algorithm 1), and we do not freeze the network weights (experiments showed that freezing the network made no impact on the stability of the learning).
@@
0:  population of players $P$ of size $N$, total episodes $T=30000$
1:  set random initial $environment\_state$ (i.e., one random move for each player)
2:  for all players $i\in P$ do
@@
3:     initialize selection memory $D_{i,sel}$ and dilemma memory $D_{i,dil}$ (capacity $256$ to allow experience replay)
4:     initialize value functions $Q_{i,sel},Q_{i,dil}$
5:     set random initial selection & dilemma states $(s_{i,sel}^{t=0},s_{i,dil}^{t=0})$
6:  end for
7:  initialize iterations counter $t=0$
8:  for all ${episode}\leftarrow 1,T$ do
9:     Get current global $environment\_state$ from the environment
10:     PART 1: Partner Selection
11:     for all players $i\in P$ do
12:        if ${episode}=0$ then
13:           use random state stored in $s_{sel}^{t=0}$ on first episode
14:        else
15:           $i$ observes the $environment\_state$ and removes itself to parse in their own $s_{i,sel}^{t}$
16:        end if
17:        $t\leftarrow t+1$
@@
18:        using $s_{i,sel}^{t}$, $Q_{i,sel}$ and $\epsilon$-greedy policy, $i$ selects a partner $j$ (i.e., takes action $a_{i,sel}^{t}$)
19:        store pair of agents $\{i,j\}$ in this episode’s sub-population $P_{episode}$
20:     end for
21:     PART 2: Dilemma Game
22:     for all $\{i,j\}\in P_{episode}$ do
23:        if $episode=0$ then
24:           use random dilemma state stored in $s_{dil}^{t=0}$ on first episode
25:        else
@@
26:           $i$ and $j$ parse states $s_{i,dil}^{t}$, $s_{j,dil}^{t}$ from the $environment\_state$ the current opponent’s index
27:        end if
28:        using $s_{i,dil}^{t}$, $Q_{i,dil}$ and $\epsilon$-greedy policy, $i$ plays a move $a_{i,dil}^{t}$
29:        using $s_{j,dil}^{t}$, $Q_{j,dil}$ and $\epsilon$-greedy policy, $j$ simultaneously plays a move $a_{j,dil}^{t}$
30:        $i$ and $j$ receive rewards $r_{i,dil}^{t}$ $r_{j,dil}^{t}$ respectively
31:        $i$ and $j$ observe $s_{dil}^{t+1}$ based on the current opponent’s action $a_{t}$: $s_{i,dil}^{t+1}=a_{j,dil}^{t};s_{j,dil}^{t+1}=a_{i,dil}^{t}$
32:        player $i$ copies the dilemma reward to the selection reward as well: $r_{i,sel}^{t}=r_{i,dil}^{t}$
33:        in player $i$’s $D_{i,dil}$ memory, store: $(s_{i,dil}^{t},a_{i,dil}^{t},r_{i,dil}^{t},s_{i,dil}^{t+1})$
34:        in player $j$’s $D_{j,dil}$ memory, store: $(s_{j,dil}^{t},a_{j,dil}^{t},r_{j,dil}^{t},s_{j,dil}^{t+1})$
35:     end for
36:     PART 3: Update $sel$ and $dil$ networks
37:     for all players $i\in P$ do
@@
38:        update the corresponding part of $environment\_state$ with the latest moves by $i$ ($a_{i,dil}^{t}$)
@@
39:        player $i$ observes their $s_{i,sel}^{t+1}$ by copying the latest $environment\_state$ and removing self
40:        in player $i$’s $D_{i,sel}$ memory, store: $(s_{i,sel}^{t},a_{i,sel}^{t},r_{i,sel}^{t},s_{i,sel}^{t+1})$
41:        player $i$ samples one experience from $D_{i,sel}$: $(s_{i,sel}^{t},a_{i,sel}^{t},r_{i,sel}^{t},s_{i,sel}^{t+1})$ calculates loss
42:        player $i$ samples all experiences from $D_{i,dil}$: $(s_{i,dil}^{t},a_{i,dil}^{t},r_{i,dil}^{t},s_{i,dil}^{t+1})$ calculates average loss
@@
43:        player $i$ updates $Q_{i,sel}$ network with respect to time step $t$ the calculated loss
@@
44:        player $i$ updates $Q_{i,dil}$ network with respect to time step $t$ the calculated loss
45:        refresh memory buffers $D_{i,sel}$, $D_{i,dil}$, $D_{j,dil}$
46:     end for
47:  end for